Evaluate: If the sides of a square measure 9sqrt(3) inches, then find the length of the diagonal.
(Write your answer in the square root form)

Answer by blessedbliss(1)

Let x = measure of the diagonal
By Pythagorean Theorem,
x^2 = (9sqrt(3))^ 2 + (9sqrt(3))^2
x^2 = (81*3) + (81*3)
= 243 +243
= 486
x = sqrt(486)
= sqrt(81*6)
= 9 sqrt(6) inches is the measure of the diagonal of the square